# Long-Tailed 3D Detection via 2D Late Fusion

### v3

@@ 3.1 How Do We Incorporate RGB Information?
Although LiDAR offers accurate localization, LiDAR-only detectors struggle to distinguish foreground objects from background using sparse LiDAR alone. RGB images provide complementary information that is essential for identifying objects and disambiguating those that are geometrically similar in point clouds but semantically different in images (e.g., construction worker vs. police officer). Although prior works address late-fusion, they combine 3D RGB detectors with 3D LiDAR detections [5]. In contrast, we find that ensembling 2D RGB detectors with 3D LiDAR detectors yields significantly better LT3D performance. We present insights on why using 2D detectors to incorporate RGB information yields better performance below, and ablate the impact of using 2D versus 3D RGB detectors for late-fusion in Table 2.
@@
2D RGB detectors are more mature. 2D object detection is a fundamental problem in computer vision [49, 50, 51] that has matured in recent years and model trade-offs are well understood [51, 12, 52, 53]. In this work, we consider two state-of-the-art 2D RGB detectors, YOLOV7 [54] and DINO [55]. YOLOV7 is a real-time detector that identifies a number of training tricks that nearly doubles the inference efficiency over prior work without sacrificing performance. DINO (not to be confused with the foundational feature extractor DINO [56]) is a recent transformer-based detector that improves upon DETR [13] using denoising anchor boxes. As 2D detectors do not make 3D predictions (e.g., depth and rotation), understanding how to best leverage them in the context of long-tailed 3D detection is a key challenge. We address this in Section 3.2.
@@
2D RGB detectors can be trained with more diverse data. Training 2D RGB detectors only requires 2D bounding box annotations, which are significantly cheaper to collect than 3D cuboids used for training 3D RGB detectors [7, 23]. Since annotating 3D amodal cuboids is both expensive and non-trivial (compared to bounding-box annotations for 2D detection), datasets for monocular 3D RGB detection are considerably smaller and less diverse than their 2D detection counterparts. For example, nuScenes [2] (published in 2020) annotates 144K RGB images of 23 classes using 3D cuboids, while COCO [50] (an early 2D detection dataset published in 2014) annotates 330K images of 80 classes using 2D bounding boxes. This allows us to pretrain 2D RGB detectors on significantly larger, more diverse, publicly available datasets [57, 58, 59, 60, 61]. We demonstrate that leveraging existing 2D detection datasets helps train stronger 2D detectors “for free”, further improving LT3D performance (cf. Fig. 7).
3.2 How Do We Match Uni-Modal Detections?
@@
Finding correspondence between two sets of RGB and LiDAR uni-modal detections is an essential step in the late-fusion framework (cf. Fig. 2B). Prior work [5] matches 3D RGB and 3D LiDAR detections using center distance in the bird’s-eye-view (BEV) plane. However, precisely matching detections is difficult due to depth estimation errors from 3D RGB detectors. Instead, we opt to match 2D RGB and 3D LiDAR detections. Prior work attempts to inflate 2D detections to the 3D BEV (using LiDAR points [9]), but we find that matching inflated 2D detections in the 3D BEV suffers from many of the same limitations as 3D RGB detections. Importantly, inflated 2D RGB detections impute additional depth estimation errors and reduce overall match quality. In contrast, we match multi-modal detections by projecting 3D LiDAR detections onto the 2D image plane, avoiding additional noise due to imprecise depth estimates. We ablate the impact of matching in the 3D BEV versus on the 2D image plane in Table 2, and present our 2D matching algorithm below.
@@
Spatial matching on the 2D image plane. Using the available sensor extrinstics, we project 3D LiDAR detections onto the 2D image plane. We then use the IoU metric to determine overlap between (projected) LiDAR and 2D RGB detections. A 2D RGB detection and a (projected) 3D LiDAR detection are considered a match if their IoU is greater than a fixed threshold. Although conceptually simple, we find that it works significantly better than using center distance to match detections in the 3D BEV (cf. bottom two rows in Table 2). Spatially matching uni-modal detections using 2D IoU yields three categories of detections: matched detections, unmatched RGB detections (that do not have corresponding LiDAR detections), and unmatched LiDAR detections (that do not have corresponding RGB detections).
@@
Handling unmatched detections. We remove unmatched 2D RGB detections, positing that any unmatched RGB detections are likely to be false positives given that LiDAR detectors tend to yield high recall [5]. Further, accurately inflating 2D RGB detections to 3D remains challenging. For unmatched 3D LiDAR detections, we down-weight their detection confidence scores by $w$ (inspired by SoftNMS [62]). We optimize for $w$ by grid search on a val-set, and set $w=0.4$.
Figure 3: Three examples demonstrate how ensembling 2D RGB detections (from DINO [55]) and 3D LiDAR detections (from CenterPoint [8]) improves LT3D performance. In all examples, our method is able to correctly relabel detections which are geometrically similar (w.r.t size and shape) in LiDAR but are visually distinct in RGB, such as bus-vs-truck, adult-vs-stroller, and adult-vs-child.
3.3 How Do We Fuse Matched Detections?
@@
As illustrated by Fig. 2C, detections may match spatially but not semantically. To address this, we propose a semantic matching heuristic to better fuse LiDAR and RGB detections. Given a pair of spatially matched RGB and LiDAR detections, we consider two cases: matched detections with semantic disagreement (e.g. RGB and LiDAR predict different classes), and matched detections with semantic agreement (e.g. RGB and LiDAR predict the same class).
@@
Addressing semantic disagreement between modalities. If both modalities predict different semantic classes, we use the confidence score and class label of the RGB-based detection and the 3D box extent from the LiDAR-based detection. Intuitively, RGB detectors can predict semantics more reliably from high resolution images than LiDAR-only detectors. This helps correct misclassifications of geometrically similar, but semantically different objects produced by the 3D LiDAR detector, as shown in Fig. 3. Importantly, prior late fusion methods like CLOCS [35] only perform late-fusion on matched predictions with semantic agreement and do not fix misclassifications. We find that this is critical for improving rare class performance.
@@
If both modalities predict the same semantic class, we perform score fusion and probablistic ensembling [63] as described below. Note that the confidence scores of RGB and LiDAR detections are not directly comparable by default. Specifically, we find that LiDAR-based detectors are often under-confident because it is difficult to distinguish foreground-vs-background using spare LiDAR alone. Therefore, score calibration is crucial for fusion. Below, we explore score calibration of RGB ($x_{\text{RGB}}$) and LiDAR ($x_{\text{LiDAR}}$) detections.
@@
Score calibration. In this work, we calibrate detection confidences per model by tuning a temperature $\tau_{c}$ for the logit score of class $c$ on a val-set before applying a sigmoid transform [64, 63], i.e., sigmoid(logitc/$\tau_{c}$). Optimally tuning per-class $\tau_{c}$ is computationally expensive as it requires tuning for all classes at the same time. Instead, we choose to greedily tune each $\tau_{c}$, optimizing per-class AP on a val-set for each class progressively ordered by their cardinalities. It is worth noting that this score calibration is only performed once in training and tuned $\tau_{c}$ and $p(c)$ do not need further optimization during inference. Importantly, score calibration does not increase runtime or complexity.
@@
$p(c|x_{\text{RGB}},x_{\text{LiDAR}})$ $=p(x_{\text{RGB}},x_{\text{LiDAR}}|c)p(c)\ \ /\ \ p(x_{\text{RGB}},x_{\text{LiDAR}})$
@@
where $p(c|x_{\text{RGB}})$ and $p(c|x_{\text{LiDAR}})$ are the posteriors after calibration. Unlike balanced class distribution studied in [63], in the long-tailed scenario, $p(c)$ can significantly impact the final LT3D performance. To maximize performance, the optimal tuning practice is to jointly tune all class priors $p(c)$ altogether, which is computationally expensive. Hence, similar to score calibration, we greedily tune $p(c)$, one by one ordered by class cardinality. Similar to score calibration, probabilistic ensembling does not increase inference time or complexity.
@@ 4 Experiments
Table 1: Benchmarking various approaches for LT3D on the nuScenes dataset (measured by mAP). We denote the modalities used in each method: C for RGB camera and L for LiDAR. Recall that our approach fuses 3D LiDAR and 2D RGB detections (produced by CenterPoint and DINO, respectively) on the 2D image plane with score calibration and probabilistic ensembling. Our frustratingly simple approach performs the best averaged across all common and rare classes, notably outperforming end-to-end multi-modal methods such as BEVFusion [6], DeepInteraction [65], and CMT [66]. Note that “CenterPoint + FCOS3D” is a baseline late-fusion method proposed by [5], which underperforms our late-fusion approach. Importantly, our approach nearly doubles the detection performance achieved by prior work on classes with few examples!
@@
      CenterPoint + FCOS3D [5]       C+L       43.6       77.1       49.0       9.4
      CenterPoint + DINO Det. (Ours)       C+L       51.4       77.9       59.4       20.0
In this section, we conduct extensive experiments to evaluate our proposed approach. We compare our late-fusion approach with prior work and present a detailed ablation study that further addresses the three motivating questions in Fig. 2. We find that our proposed approach improves over prior works by 5.9% mAP, notably improving by 7.2% on rare classes (Table 1). We include our code in the supplementary material, and will publically release it to facilitate future work.
Datasets. We study LT3D using the well-established nuScenes dataset [2]. We follow the suggested protocol in [5], using all 18 long-tail categories for benchmarking. Moreover, we use the nuImages dataset [2] as an external 2D annotated data source to study how using additional data to train better 2D RGB detectors improves late-fusion performance. Lastly, we also evaluate our approach on the Argoverse 2 dataset [3] in the supplement, and find that our conclusions hold across datasets. Note that other datasets such as KITTI [1] and Waymo [4] do not support the study of LT3D as they only annotate three common classes.
Metrics. Mean average precision (mAP) is an established metric for object detection [50]. For 3D detection, a true positive (TP) is defined as a detection that has a center distance within a distance threshold on the ground-plane to a ground-truth annotation [2]. mAP computes the mean of AP over all classes, where per-class AP is the area under the precision-recall curve drawn with distance thresholds of [0.5, 1, 2, 4] meters. We report the metrics for three groups of classes based on their cardinalities: Many ($>$50k training instances per class), Medium (5$\sim$50k), and Few ($<$5k). Following [5], we use the official nuScenes train-set for training and report numbers on the val-set because the official test-set does not evaluate on all 18 LT3D classes.
4.1 Comparison to the State-of-the-Art Methods
@@
We compare our late-fusion approach against prior work in Table 1, and present qualitative results in Fig. 3. We adapt existing methods (which were previously trained on 10 classes in nuScenes) for LT3D by retraining them on all 18 classes.
@@
CenterPoint [8], a popular LiDAR-only 3D detector, is unable to detect rare objects, achieving just 3.5% AP on classes with few examples. This is expected as it is difficult to identify rare objects from sparse LiDAR points alone. Perhaps surprisingly, the transformer-based 3D LiDAR detector BEVFusion-L performs considerably better on rare classes, achieving 10.6% AP. However, BEVFusion-L performs 3.9% worse than CenterPoint on classes with many examples. We posit that limited training data in-the-tail and class imbalance makes it difficult to learn generalizable features. In contrast, BEVFusion [21], which is an end-to-end trained multi-modal method, performs 3.0% better than the LiDAR-only variant (BEVFusion-L), confirming the benefit of using both RGB and LiDAR for LT3D. Next, we implement a simple baseline (termed “CenterPoint + RCNN”) that trains a region-based CNN (RCNN) [67] classifier on cropped regions corresponding to projected 3D detections. Notably, it underperforms CenterPoint by 6.4%, presumably because learning classifiers on cropped regions does not exploit contextual information and leads to worse classification accuracy. Peri et. al. [5] introduces a simple filtering algorithm (termed “CenterPoint + FCOS3D”), which keeps CenterPoint detections that are close (based on center distance) to monocular 3D RGB detections produced by FCOS3D in the BEV and discards all other LiDAR predictions. Given the success of this simple approach, we study this late-fusion paradigm further. Lastly, we re-implement CLOCS [35] by fusing detections from DINO and CenterPoint respectively. Note that CLOCS only fuses predictions from the same class, which prevents re-labeling misclassified LiDAR detections.

### v4

@@ 4.4.1 How Do We Incorporate RGB Information?
Although LiDAR offers accurate localization, LiDAR-only detectors struggle to recognize objects using sparse LiDAR alone. RGB images provide complementary information that is essential for identifying objects and disambiguating those that are geometrically similar in point clouds but semantically different in images, e.g., construction worker vs. police officer. In the last subsection, we proposed multimodal filtering (MMF) (Fig. 5), a strategy that ensembles a 3D RGB detector and a 3D LiDAR detector, yielding remarkable improvements for LT3D (Table II). However, we find that ensembling a 2D RGB detector and a 3D LiDAR detector yields more significantly better LT3D performance. We present insights on why using a 2D detector to incorporate RGB information is better and ablate the impact of using 2D vs. 3D RGB detectors for late-fusion in Table III.
@@
2D RGB detectors are more mature. 2D object detection is a fundamental problem in computer vision [62, 55, 63] that has matured in recent years and model trade-offs are well understood [63, 15, 64, 60]. In this work, we consider two state-of-the-art 2D RGB detectors, YOLOV7 [58] and DINO [11]. YOLOV7 is a real-time detector that identifies a number of training techniques that nearly doubles the inference efficiency over prior work without sacrificing performance. DINO [11] is a recent transformer-based detector that improves upon DETR [16] using denoising anchor boxes. As 2D detectors do not make 3D predictions (e.g., depth and rotation), understanding how to best leverage them in the context of LT3D is a key challenge.
@@
2D RGB detectors can be trained with more diverse data. Training 2D RGB detectors only requires 2D bounding box annotations, which are significantly cheaper to collect than 3D cuboids used for training 3D RGB detectors [13, 25]. Since annotating 3D amodal cuboids is both expensive and non-trivial (compared to bounding-box annotations for 2D detection), datasets for monocular 3D RGB detection are considerably smaller and less diverse than their 2D detection counterparts. For example, nuScenes [6] (published in 2020) annotates 144K RGB images of 23 classes using 3D cuboids, while COCO [55] (an early 2D detection dataset published in 2014) annotates 330K images of 80 classes using 2D bounding boxes. This allows us to pretrain 2D RGB detectors on significantly larger, more diverse, publicly available datasets [65, 66, 67, 68, 69]. We demonstrate in Fig. 11 that leveraging existing 2D detection datasets helps train stronger 2D detectors, further improving LT3D performance.
Figure 8: Our multi-modal late-fusion (MMLF) strategy takes 3D LiDAR and 2D RGB detections as input (Fig. 3A), matches 2D RGB and (projected) 3D LiDAR detections on the image plane (Fig. 3B), and fuses matched predictions with score calibration and probabilistic ensembling to produce 3D detections (Fig. 3C).
4.4.2 How Do We Match Uni-Modal Detections?
@@
Finding correspondence between two sets of RGB and LiDAR uni-modal detections is an essential step in the late-fusion framework (Fig. 3B). Our previous multi-modal filtering method [9] matched 3D RGB and 3D LiDAR detections using center distance (Fig. 5). However, precisely matching detections in 3D is difficult due to depth estimation errors from 3D RGB detectors. Instead, we opt to match 2D RGB and 3D LiDAR detections. Prior work attempts to inflate 2D detections to 3D (using LiDAR points [70]), but we find that matching inflated 2D detections also yields noisy 3D boxes (due to outlier points from the background) and reduces overall match quality. In contrast, we match uni-modal detections by projecting 3D LiDAR detections onto the 2D image plane, avoiding additional noise due to imprecise depth estimates. We ablate the impact of matching in 3D versus on the 2D image plane in Table III, and present our 2D matching algorithm below.
@@
Spatial matching on the 2D image plane. Using the available sensor extrinstics, we project 3D LiDAR detections onto the 2D image plane. We then use the IoU metric to determine overlap between (projected) LiDAR and 2D RGB detections. A 2D RGB detection and a (projected) 3D LiDAR detection are considered a match if their IoU is greater than a threshold. Although conceptually simple, this matching method works significantly better than using center distance to match detections in 3D (cf. bottom two rows in Table III). Spatially matching uni-modal detections using 2D IoU yields three categories of detections: matched detections, unmatched RGB detections (that do not have corresponding LiDAR detections), and unmatched LiDAR detections (that do not have corresponding RGB detections). We handle unmatched detections below and present our fusion strategy for matched detections in the next subsection.
@@
Handling unmatched detections. We remove unmatched 2D RGB detections, positing that any unmatched RGB detections are likely to be false positives given that LiDAR detectors tend to yield high recall [9]. Further, accurately inflating 2D RGB detections to 3D remains challenging. For unmatched 3D LiDAR detections, we down-weight their detection confidence scores by $w$ (inspired by SoftNMS [71]). We optimize for $w$ by grid search on a validation set and set $w=0.4$.
4.4.3 How Do We Fuse Matched Uni-Modal Detections?
@@
As illustrated by Fig. 3C and Fig. 9, detections may match spatially but not semantically. To address this, we propose a semantic matching heuristic. Given a pair of spatially matched RGB and LiDAR detections, we consider two cases: matched detections with semantic disagreement (e.g., RGB and LiDAR predict different classes), and matched detections with semantic agreement (e.g., RGB and LiDAR predict the same class).
@@
Addressing semantic disagreement between modalities. If the two modalities predict different semantic classes, we use the confidence score (which is calibrated as explained below) and class label of the RGB-based detection and the 3D box extent from the LiDAR-based detection. Intuitively, RGB detectors can predict semantics more reliably from high resolution images than LiDAR-only detectors. This helps correct misclassifications of geometrically similar but semantically different objects produced by the 3D LiDAR detector, as shown in Fig. 9. Importantly, prior late-fusion methods like CLOCS [38] and our MMF only perform late-fusion on matched predictions with semantic agreement and do not fix misclassifications. We find that handling such misclassifications is critical for improving rare class performance (Table I).
@@
If both modalities predict the same semantic class, we perform score fusion and probablistic ensembling [72] as described below. Note that the confidence scores of RGB and LiDAR detections are not directly comparable by default: LiDAR-based detectors are often under-confident as it is difficult to distinguish foreground-vs-background using sparse LiDAR alone. Therefore, score calibration is crucial for fusion. Below, we explore score calibration of RGB ($x_{\text{RGB}}$) and LiDAR ($x_{\text{LiDAR}}$) detections.
@@
Score calibration. We calibrate detection confidences per model by tuning a temperature $\tau_{c}$ for the logit score of class $c$ on a validation set before applying a sigmoid transform [73, 72], i.e., sigmoid(logitc/$\tau_{c}$). Optimally tuning per-class $\tau_{c}$ is computationally expensive as it requires tuning for all classes at the same time. Instead, we choose to greedily tune each $\tau_{c}$, optimizing per-class AP on a val-set for each class progressively ordered by their cardinalities. It is worth noting that this score calibration is only performed once in training and tuned $\tau_{c}$ and $p(c)$ do not need further optimization during inference. Importantly, score calibration does not increase runtime or complexity.
@@
$\small p(c|x_{\text{RGB}},x_{\text{LiDAR}})$ $=p(x_{\text{RGB}},x_{\text{LiDAR}}|c)p(c)\ \ /\ \ p(x_{\text{RGB}},x_{\text{LiDAR}})$
@@
where $p(c|x_{\text{RGB}})$ and $p(c|x_{\text{LiDAR}})$ are the posteriors after calibration. Unlike the balanced class distribution studied in [72], $p(c)$ can significantly impact the final long-tailed performance. To maximize detection accuracy, the optimal tuning practice is to jointly optimize all class priors $p(c)$ altogether, which is computationally expensive. Similar to score calibration, we greedily tune $p(c)$, one by one ordered by class cardinality. Probabilistic ensembling does not increase inference time or complexity.
@@ 5 Experiments
We conduct extensive experiments to better understand the LT3D  problem and gain insights by validating our techniques described in Sec. 4. Specifically, we aim to answer the following questions: [^1] Answers: yes, yes, yes, yes.
  1. 1.
    Are rare classes more difficult to detect than common classes?
  2. 2.
    Are objects from rare classes sufficiently localized but mis-classified?
  3. 3.
    Does training with the semantic hierarchy improve detection performance for LT3D?
  4. 4.
    Does multi-modal late-fusion help detect rare classes?
We compare our multi-modal late-fusion (MMLF) approaches (Fig. 5 and 8) with prior works and present a detailed ablation study to further addresses the three motivating questions in Fig. 3. Our approach improves over prior works by 5.9 mAP, notably improving by 7.2 mAP on rare classes (Table I).
Datasets. To explore LT3D, we use the well established nuScenes [6] and Argoverse 2.0 (AV2) [4] datasets. Other AV datasets such as KITTI [2] and Waymo [5] do not support the study of LT3D as they only annotate three common classes. nuScenes and AV2 have 18 and 26 fine-grained classes, respectively, which follow long-tailed distributions. To quantify the long-tail, we compute the imbalance factor (IF), defined as the ratio between the numbers of annotations of the most and the least common classes [44]: nuScenes and AV2 have an IF=1,670 and 2,500 respectively, which are significantly more imbalanced than existing long-tail image recognition benchmarks, e.g., iNaturalist (IF=500) [74] and ImageNet-LT (IF=1,000) [75]. nuScenes arranges classes in a semantic hierarchy (Fig. 4); AV2 does not provide a semantic hierarchy but we construct one based on the nuScenes’ hierarchy (please refer to the supplement for details). Following prior work, we use official training and validation sets. We focus on nuScenes in the main paper and AV2 in the supplement. Our primary conclusions hold for both datasets.
Figure 9: Three examples demonstrate how our multi-modal late-fusion (MMLF) approach improves LT3D by ensembling 2D RGB detections (from DINO [11]) and 3D LiDAR detections (from CenterPoint [12]). In all examples, MMLF correctly relabels detections which are geometrically similar (w.r.t size and shape) in LiDAR but are visually distinct in RGB, such as bus-vs-truck, adult-vs-stroller, and adult-vs-child.
TABLE I: Benchmarking results on nuScenes. We denote the RGB and LiDAR modalities by C and L, respectively. Our MMLF fuses 3D LiDAR and 2D RGB detections (from CenterPoint and DINO, respectively) with score calibration and probabilistic ensembling. MMLF performs the best averaged across all common and rare classes, notably outperforming end-to-end multi-modal methods such as BEVFusion [35], DeepInteraction [59], and CMT [10]. Importantly, our MMLF approach nearly doubles the detection performance achieved by prior work on classes with few examples!
@@
CenterPoint [12] + RCNN [76] C+L 34.0 64.8 37.5 4.3
MMF(CenterPoint, FCOS3D) (Ours) C+L 43.6 77.1 49.0 9.4
MMLF(CenterPoint, DINO) (Ours) C+L 51.4 77.9 59.4 20.0
5.1 State-of-the-Art Comparison on nuScenes
@@
Table I compares our MMF and MMLF approaches with prior work on nuScenes. Fig. 9 presents qualitative results. We adapt existing methods (which were previously trained on 10 classes in nuScenes) for LT3D by retraining them on all 18 classes.
@@
CenterPoint [12], a popular LiDAR-only 3D detector, is unable to detect rare objects, achieving just 4.3 mAP on classes with few examples. This is expected as it is difficult to identify rare objects from sparse LiDAR points alone. Perhaps surprisingly, the transformer-based 3D LiDAR detector BEVFusion-L performs considerably better on few classes, achieving 10.6 mAP, but it underperforms CenterPoint by 4.6 mAP on classes with many examples. We posit that limited training data in-the-tail and class imbalance make it difficult to learn generalizable features, preventing robust LT3D performance. In contrast, BEVFusion [23], which is an end-to-end trained multi-modal method, performs 3.0 mAP better than the LiDAR-only variant (BEVFusion-L), confirming the benefit of using both RGB and LiDAR for LT3D.
Next, we implement another simple baseline, termed “CenterPoint + RCNN”, that trains a region-based CNN (RCNN) [76] classifier on cropped regions corresponding to projected 3D detections. Notably, it underperforms CenterPoint by 6.4 mAP, presumably because learning classifiers on cropped regions does not exploit contextual information and leads to worse classification accuracy. This suggests that late-fusion cannot be simply solved with a 3D LiDAR detector and a strong 2D classifier. Our multimodal filtering method MMF(CenterPoint, FCOS3D) keeps CenterPoint detections that are close to monocular 3D RGB detections produced by FCOS3D in 3D and discards all other LiDAR predictions. This simple baseline achieves 9.4 mAP on classes with few examples. Lastly, we re-implement CLOCS [38] by fusing detections from DINO and CenterPoint respectively. Note that CLOCS only fuses predictions from the same class, which prevents re-labeling misclassified LiDAR detections. However, CLOCs performs worse than our MMF baseline. By carefully considering design choices outlined in Fig. 3, MMLF(CenterPoint, DINO) improves over MMF(CenterPoint, FCOS3D) by 7.8 mAP!
5.2 Ablation Studies and Empirical Analysis
We conduct extensive ablations to validate our proposed MMLF approach. We provide additional results in the supplement.
TABLE II: Ablation on Hierarchy Loss and Multi-Modal Filtering (MMF). Training with the semantic hierarchy improves both PointPillars [17] and CenterPoint [12] for LT3D by $>$1% mAP over All classes. Moreover, MMF yields 4$\sim$11 mAP improvement on Medium and Few classes, showing the benefit of late fusion!
Method Modality All Many Medium Few
FCOS3D [13] C 20.9 39.0 23.3 2.9
PointPillars [17] L 30.0 64.2 28.4 3.4
+ Hierarchy Loss L 31.2 66.4 30.4 2.9
+ MMF(PointPillars, FCOS3D) C+L 35.8 66.2 41.0 4.4
CenterPoint[12] L 39.2 76.4 43.1 3.5
+ Hierarchy Loss L 40.4 77.1 45.1 4.3
+ MMF(CenterPoint, FCOS3D) C+L 43.6 77.1 49.0 9.4
5.2.1 Analysis on Training with Semantic Hierarchy
We validate the benefits of training with semantic hierarchy by modifying LiDAR-based detectors (PointPillars [17] and CenterPoint [12]) to jointly predict class labels at different levels of the semantic hierarchy. For example, we modify the detector to additionally classify stroller as pedestrian and object. The semantic hierarchy naturally groups classes based on shared attributes and may have complementary features, helping train a detector that allows feature sharing across common and rare classes and leads to improved performance. Table II highlights that this approach (“+ Hierarchy Loss”) effectively improves LT3D accuracy by $>$1 mAP over all classes.